OpenStudy (anonymous):

A flower bed is in the shape of a triangle with one side twice the length of the shortest side, and the third side is 13 feet more than the length of the shortest side. Find the dimensions if the perimeter is 125 feet

OpenStudy (across):



OpenStudy (anonymous):

perimeter is sum of all sides

OpenStudy (across):

kylemp1, does that drawing follow the description?

OpenStudy (anonymous):

therefore 2x + x + x + 13 = 125 4x = 125-13 4x = 112 x = 28

OpenStudy (anonymous):

I'm pretty sure, it just gives me the question, but no drawing

OpenStudy (anonymous):

there fore dimension is x = 28 2x = 56 x+13 = 41

OpenStudy (anonymous):

so the shortest side is 28, the second side is 41, and the third is side is 51?

OpenStudy (across):

Look at the drawing and solve the equation\[(x)+(2x)+(x+13)=125.\]

OpenStudy (across):

Then check your answer with that of helpingtutors.